# NS-Net: Decoupling CLIP Semantic Information through NULL-Space for Generalizable AI-Generated Image Detection

### v1

@@ Introduction
Analogously, our primary objective is to construct the NULL-Space corresponding to the semantic information embedded within the visual features extracted by CLIP’s image encoder. Specifically, we use the CLIP’s image encoder to extract visual features $U$ from each image, which inherently encode both semantic information $U_{s}$ and artifact-related information $U_{t}$, as discussed in Section I. However, the entanglement and complex interactions between $U_{s}$ and $U_{t}$ hinder the direct construction of the NULL-space $X_{U_{s}}$ associated with semantic content alone. Given that CLIP is pre-trained for image-text alignment, the textual features, extracted by CLIP’s text encoder from the corresponding image description, should ideally align with the semantic component $U_{s}$, i.e., $V\approx U_{s}$. Consequently, we shift our goal to constructing the NULL-Space $X_{V}$ of the textual features $V$, and then projecting the visual features $U$ onto $X_{V}$ in order to suppress semantic content and achieve effective disentanglement. During training, we freeze both CLIP’s image and text encoders and introduce LoRA (Low-Rank Adaptation) layers to fine-tune the image encoder. This design improves the encoder’s ability to emphasize artifact-related cues relevant for distinguishing AI-generated images.
To begin, we consider a dataset $D$ consisting of both real and fake images, defined as:
$D=\{\text{img}_{j},\text{label}_{j}\}_{j=1}^{N},\text{label}\in\{0,1\},$ (2)
[…]
$\tilde{D}=\{\text{img}_{j},\text{text}_{j},\text{label}_{j}\}_{j=1}^{N},\text{label}\in\{0,1\}.$ (3)
@@
[…]
Specifically, we first apply Singular Value Decomposition (SVD) (Wang et al. 2021) to $V$:
@@
Contrastive Learning and Classification
We believe that direct binary classification will cause the network to overfit to a specific feature structure, which will inhibit the generalization effect of the detector. In order to make the detector more focused on learning the difference between the features of real and generated images, we use contrastive loss (Chen et al. 2020) $\mathcal{L}_{\text{contrastive}}$ as the dominant and linear classifier loss $\mathcal{L}_{\text{BCE}}$ as the auxiliary for training.
Specifically, for the decoupled features, we define features with the same label as positive samples and features with different labels as negative. We hope that positive samples are closer, while negative samples are gradually farther away. Thus, let $i\in I\equiv\{1...N\}$ be the index of the batch sample, where N is the batchsize. The self-supervised contrastive loss of decoupled features can be formulated as follow:
where $f$ are the decoupled features, the index $i$ called the anchor, the index $p$ called the positive, $\mathcal{P}_{i}=\{j\mid\text{label}_{i}=\text{label}_{j},\,j\neq i\}$ is the set of positive samples, $\tau$ is a temperature hyperparameter and ||$\cdot$|| ⋅ | denotes the number of vectors.
where $\lambda$ is the hyper-parameter for balancing two losses.
@@
Generative Adversarial Networks Other
ProGAN
CycleGAN
BigGAN
StyleGAN
StyleGAN2
GauGAN
StarGAN
@@
CNN-Spot 99.0 95.3 18.7 1.8 36.5 22.0 2.5 23.1 1.1
UnivFD 92.3 98.9 90.5 79.2 55.7 48.7 91.1 96.9 92.8
FreqNet 89.9 99.4 57.1 51.0 75.1 67.5 9.9 88.4 95.4
NPR 99.3 98.9 29.3 16.5 67.1 58.7 1.7 6.5 7.9
Ladeda 99.8 99.7 7.2 29.0 95.6 98.3 4.9 0.0 19.2
AIDE 93.0 95.3 88.0 96.9 89.6 97.0 90.0 97.0 42.9
DFFreq 98.0 98.0 70.9 93.6 99.0 98.9 93.8 97.4 3.5
SAFE 99.2 99.7 85.6 83.5 88.1 96.7 89.0 99.9 8.4
VIB-Net 93.5 99.4 97.3 91.3 71.3 73.3 97.8 97.3 97.3
Ours 97.7 100.0 99.9 100.0 95.7 98.6 98.5 100.0 68.9
Method Other Diffusion Models mAcc.
Deepfake
@@
CNN-Spot 29.3 55.9 55.6 1.8 4.8 5.2 27.6 0.7 4.5 29.0
UnivFD 26.9 96.3 96.0 12.7 75.6 61.2 84.7 45.6 62.3 72.7
FreqNet 35.8 99.9 99.8 37.7 78.9 80.8 98.0 34.1 88.8 71.7
NPR 0.1 100.0 99.9 26.5 69.2 71.0 97.7 15.4 89.8 53.1
Ladeda 0.1 100.0 99.9 27.3 79.7 88.8 97.9 15.5 92.4 58.6
AIDE 9.2 99.8 99.7 92.4 98.7 63.7 98.9 90.1 98.9 85.6
DFFreq 76.0 99.9 99.8 65.9 87.8 94.8 99.1 76.0 95.9 85.2
SAFE 17.3 99.8 99.6 36.8 90.5 86.3 98.5 84.0 92.0 80.8
VIB-Net 90.6 100.0 99.8 52.8 69.3 63.7 99.3 80.9 58.5 84.7
Ours 60.6 100.0 99.9 85.3 97.2 77.4 100.0 98.9 98.8 93.2
@@
Table 1: Cross-model Accuracy (Acc.) Performance on the GenImage (Zhu et al. 2024) Dataset. The first column represents the accuracy of detecting real images (R.Acc.), and the others are the accuracy of detecting fake images (F.Acc.).
Experiments
Evaluation Setup
Training datasets. We utilize the dataset from AIGIBench (Li et al. 2025d), specifically adopting Setting-II, which involves training on 144k images generated by ProGAN (Karras et al. 2018) and SDv1.4 (Rombach et al. 2022). These images span four object categories: car, cat, chair, and horse. In contrast to prior approaches that trained on images from a single generation technique, such as ProGAN or SDv1.4, our method incorporates multiple known generation techniques into the training dataset, aiming to improve the model’s generalization across various generation methods.
Testing datasets. To comprehensively evaluate the effectiveness of our method, we conduct generalization experiments on three different datasets, including GenImage (Zhu et al. 2024) (with deepfakes from the ForenSynths (Wang et al. 2020) dataset added), UniversalFakeDetect (Ojha, Li, and Lee 2023), and AIGIBench (Li et al. 2025d). The test set comprises 19 subsets from various generative models, including: ProGAN (Karras et al. 2018), StyleGAN, StyleGAN2, StyleGAN3 (Karras et al. 2021), StyleGAN-XL (Sauer, Schwarz, and Geiger 2022), StyleSwim (Zhang et al. 2022), BigGAN (Brock 2018), CycleGAN (Zhu et al. 2017), StarGAN (Choi et al. 2018), GauGAN (Park et al. 2019), R3GAN (Huang et al. 2024), Deepfake (Rossler et al. 2019), WFIR, LDM (Rombach et al. 2022), Midjourney, SDv1.4 (Rombach et al. 2022), SDv1.5, SD-3, SD-XL, ADM (Dhariwal and Nichol 2021), GLIDE (Nichol et al. 2021), Wukong, VQDM, Midjourney-v6, FLUX.1-dev, Imagen-3, Infinite-ID (Wu et al. 2024b), BLIP (Li et al. 2022), IP-Adapter (Ye et al. 2023), and 2 general subsets from open-source platforms, which called CommunityAI and SocialRF.
Evaluation Metrics. Following the established evaluation paradigm (Li et al. 2025d), we adopt the accuracy of detecting fake images (F.Acc.) as the primary evaluation metrics. Additionally, we calculate the accuracy of detecting real images (R.Acc.) at the beginning of the table to provide a complete assessment of the model’s performance. Additional metric about A.P. is shown in Appendix II.
Comparison Baselines. We compare our method with existing state-of-the-art detection methods, including CNN-Spot (CVPR 2020) (Wang et al. 2020), UnivFD (CVPR 2023) (Ojha, Li, and Lee 2023), FreqNet (AAAI 2024) (Tan et al. 2024a), NPR (CVPR 2024) (Tan et al. 2024b), Ladeda (arXiv 2024) (Cavia et al. 2024), AIDE (ICLR 2025) (Yan et al. 2024), DFFreq (arXiv 2025) (Yan et al. 2025), SAFE (KDD 2025) (Li et al. 2025b), and VIB-Net (CVPR 2025) (Zhang et al. 2025). More details are shown in Appendix I. To ensure a fair comparison, we retrain all methods on the AIGIBench training set.
Implementation Details. We use the Adam optimizer (Kingma 2014) with a learning rate of $2\times 10^{-4}$. The batchsize is set to 32, and we train the model for only 2 epoches. The ViT-L/14 model of CLIP is adopted as the pre-trained model following the baseline UnivFD (Ojha, Li, and Lee 2023). In addition, we add Low-Rank Adaptation to fine-tune the CLIP’s image encoder, following (Zanella and Ben Ayed 2024). For the setting of the LoRA layers, we configure the hyperparameters as follows: lora_r = 6, lora_alpha = 6, and lora_dropout = 0.8. The hyperparameters $\lambda$ is set to 0.2. The proposed method is implemented using PyTorch (Paszke et al. 2019) on an NVIDIA RTX 4090 GPU.
Method Guided Glide_50_27 Glide_100_10 Glide_100_27 LDM_100 LDM_200 LDM_200_cfg DALLE mAcc.
CNN-Spot 3.0 4.6 4.6 5.1 21.4 24.1 36.1 3.5 12.8
UnivFD 35.5 79.1 80.0 79.6 75.9 76.0 57.7 49.1 66.4
FreqNet 40.9 80.4 81.1 78.2 96.6 97.6 96.1 40.1 76.4
NPR 29.5 61.9 60.6 59.9 76.1 73.9 89.3 27.3 59.8
Ladeda 39.5 69.4 71.6 72.4 74.3 75.7 85.3 9.2 62.2
AIDE 81.7 97.3 96.5 97.0 98.6 98.4 99.7 96.1 95.6
DFFreq 80.7 90.7 92.5 88.6 99.9 99.8 99.7 91.4 92.9
SAFE 45.5 92.9 94.7 90.2 99.8 99.9 99.7 96.1 89.9
VIB-Net 58.1 75.5 76.6 72.4 97.1 97.0 86.5 93.9 82.1
Ours 85.3 96.7 98.0 96.4 100.0 100.0 100.0 100.0 97.0
Table 2: Cross-model Fake Accuracy(F.Acc.) Performance on the UniversalFakeDetect (Ojha, Li, and Lee 2023) Dataset.
@@
Generator CNN-Spot UnivFD FreqNet NPR Ladeda AIDE DFFreq SAFE VIB-Net Ours
@@
Table 3: Cross-model Accuracy (Acc.) Performance on the AIGIBench (Li et al. 2025d) Dataset. The first row represents the accuracy of detecting real images (R.Acc.), and the others are the accuracy of detecting fake images (F.Acc.).
Comparison with the State-of-the-Art

### v2

@@ 3.2 NULL-Space Decoupling
Analogously, our primary objective is to construct the NULL-Space corresponding to the semantic information embedded within the visual features extracted by CLIP’s image encoder. Specifically, we use the CLIP’s image encoder to extract visual features $U$ from each image, which inherently encode both semantic information $U_{s}$ and artifact-related information $U_{t}$, as discussed in Sec.˜1. However, the entanglement and complex interactions between $U_{s}$ and $U_{t}$ hinder the direct construction of the NULL-Space $X_{U_{s}}$ associated with semantic content alone. Given that CLIP is pre-trained for image-text alignment, the textual features, extracted by CLIP’s text encoder from the corresponding image description, should ideally align with the semantic component $U_{s}$, i.e., $V\approx U_{s}$. Consequently, we shift our goal to constructing the NULL-Space $X_{V}$ of the textual features $V$, and then projecting the visual features $U$ onto $X_{V}$ in order to suppress semantic content and achieve effective disentanglement. During training, we freeze both CLIP’s image and text encoders and introduce LoRA (Low-Rank Adaptation) layers to fine-tune the image encoder. This design improves the encoder’s ability to emphasize artifact-related cues relevant for distinguishing AI-generated images.
To begin, we consider a dataset $\mathcal{D}$ consisting of both real and fake images, defined as:
$\mathcal{D}=\{\text{img}_{j},\text{label}_{j}\}_{j=1}^{N},\;\;\text{label}\in\{0,1\},$ (2)
where $\text{label}=1$ denotes a fake image, and $\text{label}=0$ denotes a real image. For each image $\text{img}_{j}$, we employ BLIP [li2022blip] to generate a corresponding text description $\text{text}_{j}$. This yields an enriched dataset with paired visual and textual information, defined as:
$\tilde{\mathcal{D}}=\{\text{img}_{j},\text{text}_{j},\text{label}_{j}\}_{j=1}^{N},\;\;\text{label}\in\{0,1\}.$ (3)
@@
$\text{NULL-Space}(V)\;=\;\{X_{V}:VX_{V}=0\}.$ (5)
Specifically, we first apply Singular Value Decomposition (SVD) [wang2021training] to $V$:
@@
Generative Adversarial Networks Other Diffusion Models mAcc.
Pro-
GAN
Cycle-
GAN
Big-
GAN
Style-
GAN
Style-
GAN2
Gau-
GAN
Star-
GAN
@@
Deep-
fake
@@
CNN-Spot [wang2020cnn] 99.0 95.3 18.7 1.8 36.5 22.0 2.5 23.1 1.1 29.3 55.9 55.6 1.8 4.8 5.2 27.6 0.7 4.5 29.0
UnivFD [ojha2023towards] 92.3 98.9 90.5 79.2 55.7 48.7 91.1 96.9 92.8 26.9 96.3 96.0 12.7 75.6 61.2 84.7 45.6 62.3 72.7
FreqNet [tan2024frequency] 89.9 99.4 57.1 51.0 75.1 67.5 9.9 88.4 95.4 35.8 99.9 99.8 37.7 78.9 80.8 98.0 34.1 88.8 71.7
NPR [tan2024rethinking] 99.3 98.9 29.3 16.5 67.1 58.7 1.7 6.5 7.9 0.1 100.0 99.9 26.5 69.2 71.0 97.7 15.4 89.8 53.1
Ladeda [cavia2024real] 99.8 99.7 7.2 29.0 95.6 98.3 4.9 0.0 19.2 0.1 100.0 99.9 27.3 79.7 88.8 97.9 15.5 92.4 58.6
AIDE [yan2024sanity] 93.0 95.3 88.0 96.9 89.6 97.0 90.0 97.0 42.9 9.2 99.8 99.7 92.4 98.7 63.7 98.9 90.1 98.9 85.6
C2P-CLIP* [tan2025c2p] 99.8 100.0 99.8 99.8 72.2 78.0 87.7 100.0 34.9 67.6 100.0 99.9 50.5 73.3 28.4 99.7 93.3 98.1 82.4
DFFreq [yan2025generalizable] 98.0 98.0 70.9 93.6 99.0 98.9 93.8 97.4 3.5 76.0 99.9 99.8 65.9 87.8 94.8 99.1 76.0 95.9 85.2
SAFE [li2025improving] 99.2 99.7 85.6 83.5 88.1 96.7 89.0 99.9 8.4 17.3 99.8 99.6 36.8 90.5 86.3 98.5 84.0 92.0 80.8
VIB-Net [zhang2025towards] 93.5 99.4 97.3 91.3 71.3 73.3 97.8 97.3 97.3 90.6 100.0 99.8 52.8 69.3 63.7 99.3 80.9 58.5 84.7
Ours 97.7 100.0 99.9 100.0 95.7 98.6 98.5 100.0 68.9 60.6 100.0 99.9 85.3 97.2 77.4 100.0 98.9 98.8 93.2
@@
Table 1: Cross-model Accuracy (Acc.) Performance on the GenImage [zhu2024genimage] Dataset. The first column represents the accuracy of detecting real images (R.Acc.), and the others are the accuracy of detecting fake images (F.Acc.).
Method Generative Adversarial Networks Other Diffusion Models mA.P.
Pro-
GAN
Cycle-
GAN
Big-
GAN
Style-
GAN
Style-
GAN2
Gau-
GAN
Star-
GAN
WFIR
Deep-
fake
SDv1.4
SDv1.5
ADM
GLIDE
Mid-
journey
Wukong
VQDM
DALLE2
CNN-Spot [wang2020cnn] 99.9 91.0 59.2 94.8 94.2 86.1 82.5 65.5 74.2 97.6 97.6 66.6 83.1 80.4 88.9 57.0 87.8 84.3
UnivFD [ojha2023towards] 99.9 90.5 79.2 55.7 48.7 91.1 96.9 92.8 26.9 96.3 96.0 65.2 95.3 91.7 97.4 87.2 62.3 91.9
FreqNet [tan2024frequency] 99.4 97.5 87.9 95.9 94.2 54.0 100.0 57.0 88.5 99.9 99.9 73.9 94.1 94.4 99.1 74.9 92.3 88.4
NPR [tan2024rethinking] 100.0 98.6 79.4 98.6 99.6 71.3 97.6 65.5 92.5 100.0 99.9 74.1 97.4 97.8 99.9 81.5 99.3 91.4
Ladeda [cavia2024real] 100.0 91.5 93.2 100.0 100.0 81.8 93.0 86.9 72.8 100.0 99.9 77.3 99.1 99.7 100.0 89.3 99.8 93.2
AIDE [yan2024sanity] 99.6 98.2 91.6 99.4 99.9 74.1 99.7 90.8 66.5 100.0 99.9 99.4 99.9 90.3 99.9 99.3 99.9 94.5
C2P-CLIP* [tan2025c2p] 100.0 100.0 99.9 99.3 99.7 98.8 100.0 99.6 98.3 100.0 100.0 96.3 98.9 90.4 100.0 99.9 100.0 98.8
DFFreq [yan2025generalizable] 99.7 99.0 98.4 100.0 100.0 98.1 99.8 89.0 87.5 100.0 99.9 98.7 99.5 99.7 100.0 99.3 99.9 98.1
SAFE [li2025improving] 100.0 99.5 97.8 99.9 100.0 96.6 100.0 73.8 91.0 100.0 100.0 98.1 99.9 99.9 100.0 99.9 100.0 97.4
VIB-Net [zhang2025towards] 99.9 99.7 97.1 93.9 98.1 99.2 99.7 95.7 83.6 100.0 99.9 91.2 96.4 94.3 99.9 97.4 94.6 96.5
Ours 100.0 100.0 99.7 99.8 99.9 97.0 100.0 99.5 95.2 100.0 100.0 100.0 99.8 97.0 100.0 100.0 100.0 99.2
Table 2: Cross-model Accuracy (A.P.) Performance on the GenImage [zhu2024genimage] Dataset.
Method Guided Glide_50_27 Glide_100_10 Glide_100_27 LDM_100 LDM_200 LDM_200_cfg DALLE Mean
F.Acc. A.P. F.Acc. A.P. F.Acc. A.P. F.Acc. A.P. F.Acc. A.P. F.Acc. A.P. F.Acc. A.P. F.Acc. A.P. R.Acc. F.Acc. Acc. A.P.
CNN-Spot [wang2020cnn] 3.0 70.8 4.6 76.9 4.6 78.0 5.1 76.8 21.4 88.7 24.1 90.1 36.1 93.8 3.5 62.7 99.8 12.8 56.3 79.7
UnivFD [ojha2023towards] 35.5 81.0 79.1 92.5 80.0 92.1 79.6 92.3 75.9 91.4 76.0 91.4 57.7 84.7 49.1 80.3 89.8 66.4 78.1 88.2
FreqNet [tan2024frequency] 40.9 78.5 80.4 97.8 81.1 97.9 78.2 97.6 96.6 99.7 97.6 99.8 96.1 99.7 40.1 93.1 97.9 76.4 87.6 95.5
NPR [tan2024rethinking] 29.5 83.0 61.9 99.3 60.6 99.0 59.9 99.3 76.1 99.9 73.9 99.9 89.3 100.0 27.3 98.9 99.9 59.8 79.8 97.4
Ladeda [cavia2024real] 39.5 89.0 69.4 99.6 71.6 99.4 72.4 99.7 74.3 99.9 75.7 99.9 85.3 99.9 9.2 96.4 100.0 62.2 81.1 98.0
AIDE [yan2024sanity] 81.7 98.5 97.3 99.9 96.5 99.8 97.0 99.8 98.6 99.9 98.4 100.0 99.7 100.0 96.1 99.5 97.4 95.6 96.5 99.6
C2P-CLIP* [tan2025c2p] 60.7 98.6 88.7 99.6 86.0 99.8 78.6 99.4 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 89.2 94.6 99.7
DFFreq [yan2025generalizable] 80.7 99.2 90.7 99.2 92.5 99.2 88.6 99.0 99.9 100.0 99.8 100.0 99.7 100.0 91.4 99.5 98.8 92.9 96.0 99.5
SAFE [li2025improving] 45.5 98.9 92.9 99.6 94.7 99.7 90.2 99.4 99.8 100.0 99.9 100.0 99.7 100.0 96.1 99.9 99.3 89.9 94.6 99.7
VIB-Net [zhang2025towards] 58.1 94.2 75.5 97.2 76.6 96.8 72.4 96.9 97.1 99.9 97.0 99.8 86.5 98.6 93.9 99.5 98.8 82.1 90.5 97.9
Ours 85.3 99.4 96.7 99.9 98.0 99.9 96.4 99.8 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.9 97.0 98.5 99.9
Table 3: Cross-Diffusion-Sources Evaluation on the Diffusion Test Set of UniversalFakeDetect [ojha2023towards].
@@
Generator CNN-Spot [wang2020cnn] UnivFD [ojha2023towards] FreqNet [tan2024frequency] NPR [tan2024rethinking] Ladeda [cavia2024real] AIDE [yan2024sanity] DFFreq [yan2025generalizable] SAFE [li2025improving] VIB-Net [zhang2025towards] Ours
@@
Table 4: Cross-model Accuracy (Acc.) Performance on the AIGIBench [li2025artificial] Dataset. The first row represents the accuracy of detecting real images (R.Acc.), and the others are the accuracy of detecting fake images (F.Acc.).
3.3 Contrastive Learning and Classification
We believe that direct binary classification will cause the network to overfit a specific feature structure, which will inhibit the generalization effect of the detector. In order to make the detector more focused on learning the difference between the features of real and generated images, we use contrastive loss [chen2020simple] $\mathcal{L}_{\text{contrastive}}$ as the dominant and linear classifier loss $\mathcal{L}_{\text{BCE}}$ as the auxiliary for training.
Specifically, for the decoupled features, we define features with the same label as positive samples and features with different labels as negative. We hope that positive samples are closer, while negative samples are gradually farther away. Thus, let $i\in I\equiv\{1...N\}$ be the index of the batch sample, where N is the batchsize. The self-supervised contrastive loss of decoupled features can be formulated as follows:
where $f$ are the decoupled features, the index $i$ called the anchor, the index $p$ called the positive, $\mathcal{P}_{i}=\{j\mid\text{label}_{i}=\text{label}_{j},\,j\neq i\}$ is the set of positive samples, $\tau$ is a temperature hyperparameter and $|\cdot|$ denotes the number of vectors.
where $\lambda$ is the hyper-parameter for balancing two losses.
4 Experiments
4.1 Evaluation Setup
Training datasets. We use the dataset from AIGIBench [li2025artificial], specifically adopting Setting-II, which involves training on 144k images generated by ProGAN [karras2018progressive] and SDv1.4 [rombach2022high].
Testing datasets. To comprehensively evaluate the effectiveness of our method, we conduct generalization experiments on three different datasets, including GenImage [zhu2024genimage] (with deepfakes from ForenSynths [wang2020cnn] added), UniversalFakeDetect [ojha2023towards], and AIGIBench [li2025artificial]. The test set comprises 19 subsets from various generative models, including: ProGAN [karras2018progressive], StyleGAN, StyleGAN2, StyleGAN3 [karras2021alias], StyleGAN-XL [sauer2022stylegan], StyleSwim [zhang2022styleswin], BigGAN [brock2018large], CycleGAN [zhu2017unpaired], StarGAN [choi2018stargan], GauGAN [park2019semantic], R3GAN [huang2024gan], Deepfake [rossler2019faceforensics++], WFIR, LDM [rombach2022high], Midjourney, SDv1.4 [rombach2022high], SDv1.5, SD-3, SD-XL, ADM [dhariwal2021diffusion], GLIDE [nichol2021glide], Wukong, VQDM, Midjourney-v6, FLUX.1-dev, Imagen-3, Infinite-ID [wu2024infinite], BLIP [li2022blip], IP-Adapter [ye2023ip], and 2 general subsets from open-source platforms, called CommunityAI and SocialRF.
Evaluation Metrics. Following the established evaluation paradigm [li2025artificial], we adopt the accuracy of detecting fake images (F.Acc.) and average precision (A.P.) as primary evaluation metrics. Additionally, we calculate the accuracy of detecting real images (R.Acc.) at the beginning of the table to provide a complete assessment of the model’s performance.
Implementation Details. We use the Adam optimizer [kingma2014adam] with a learning rate of $1\times 10^{-4}$. The batchsize is set to 32, and we train the model for only 2 epochs. The ViT-L/14 model of CLIP is adopted as the pre-trained model following the baseline UnivFD [ojha2023towards]. In addition, we add Low-Rank Adaptation to fine-tune CLIP’s image encoder, following [zanella2024low]. For the setting of the LoRA layers, we configure the hyperparameters as follows: lora_r = 6, lora_alpha = 6, and lora_dropout = 0.8. The hyperparameter $\lambda$ is set to 0.2. The proposed method is implemented using PyTorch [paszke2019pytorch] on an NVIDIA RTX 4090 GPU. What’s more, we compare our method with 10 existing state-of-the-art detection methods. More details are shown in Sec.˜6 of the Appendix. To ensure a fair comparison, we retrain all methods on the AIGIBench.
4.2 Comparison with the State-of-the-Art